OpenStudy (anonymous):

Can someone show me how to differentiate y=6 lnx? I'm sure it's chain rule, but I'm struggling with how to set it up.

OpenStudy (mathstudent55):

\(\dfrac{d}{dx} \ln u = \dfrac{1}{u} \dfrac{d}{dx}u \)

OpenStudy (mathstudent55):

Also, \( \dfrac {d}{dx} ku = k \dfrac{d}{dx}u \)

OpenStudy (mathstudent55):

\(y = 6 \ln x\) \(y' = 6 \dfrac{d}{dx} \ln x\) \(y' = 6 \cdot\dfrac{1}{x} \cdot \dfrac{d}{dx}x\) \(y' = 6 \cdot\dfrac{1}{x} \cdot 1\) \(y' = \dfrac{6}{x}\)